# MaskInversion: Localized Embeddings via Optimization of Explainability Maps

### v1

@@ Abstract
Vision-language foundation models such as CLIP have achieved tremendous results in global vision-language alignment, but still show some limitations in creating representations for specific image regions. To address this problem, we propose MaskInversion, a method that leverages the feature representations of pre-trained foundation models, such as CLIP, to generate a context-aware embedding for a query image region specified by a mask at test time. MaskInversion starts with initializing an embedding token and compares its explainability map, derived from the foundation model, to the query mask. The embedding token is then subsequently refined to approximate the query region by minimizing the discrepancy between its explainability map and the query mask. During this process, only the embedding vector is updated, while the underlying foundation model is kept frozen allowing to use MaskInversion with any pre-trained model. As deriving the explainability map involves computing its gradient, which can be expensive, we propose a gradient decomposition strategy that simplifies this computation. The learned region representation can be used for a broad range of tasks, including open-vocabulary class retrieval, referring expression comprehension, as well as for localized captioning and image generation. We evaluate the proposed method on all those tasks on several datasets such as PascalVOC, MSCOCO, RefCOCO, and OpenImagesV7 and show its capabilities compared to other SOTA approaches. [^1] Project page: https://walidbousselham.com/MaskInversion
Figure 1: MaskInversion Applications: The proposed MaskInversion method generates a localized embedding without modifying the vision encoder, thereby enabling seamless integration as a drop-in replacement for the vision encoder output across various scenarios. (Localized Classification): classify each region of an image independently. (Localized Captioning): direct the attention of an LLM to specific parts of an image. (Localized Diffusion): used in conjunction with a diffusion model, generates variations of specific regions of images. All applications demonstrated here can be achieved by simply replacing the original vision encoder with MaskInversion, without further tuning.
@@ 1 Introduction
Foundation models such as CLIP Radford et al. (2021), pre-trained with a contrastive loss on large-scale image-text datasets, have significantly advanced vision-language understanding. However, those models focus on a global vision-language alignment in training, matching the respective text and image class ([CLS]) tokens, thus only the globally pooled information. As a result, such models often struggle with tasks requiring precise localization or the recognition of specific image regions, necessitating novel approaches to harness their full potential. In the following, we tackle the problem of generating embeddings localized to specific image regions from pretrained vision-language models. While it is possible to obtain such embeddings via naïve solutions, e.g. processing only the cropped region, or aggregating the local token embeddings over a mask, such simple approaches often do not yield optimal results: cropping can remove important context, while token aggregation over region features might not result in a good, aligned representation as local tokens do not always correspond to the correct representation Zhou et al. (2022).
@@
Different approaches have been proposed to address the problem of localized vision-language tasks: ReCLIP Subramanian et al. (2022) uses colored boxes during training to localize the alignment between vision and language. Fine-grained visual Prompting (FGVP) employs different masking strategies to force the model to focus on the relevant object region. AlphaCLIP Sun et al. (2024) finetunes CLIP together with an alpha channel to highlight the region of interest. Finally, RIS Yu et al. (2023) proposes a token masking pipeline to achieve zero-shot referring image segmentation.
Following this idea, we propose MaskInversion, inspired by Text Inversion Gal et al. (2023), as a method to learn a localized embedding for a query image region specified by a mask at test time. MaskInversion differs from previous methods as it does not adapt the vision-language backbone, but instead leverages the explainabilty map of a frozen backbone at test-time to optimize a representation, namely a token that captures the localized embedding (localized embedding token), for a given region mask. We start with initializing the localized embedding token from the global class token produced by CLIP. This token representation is then used to compute the initial explainability map for its current representation. We then compute the difference between the explainability map and the query mask. The token representation is then subsequently updated so that its representation generates an explainability map that matches the query mask. In this manner, we learn a token representation specific to the image region covered by the query mask.
@@
Note that the token representation learning process is done for each mask separately. Thus, several different localized embedding tokens are created from the same image when multiple object masks are given. We can further enhance the computational efficiency for this case by exploiting the fact that the derivation of the explainability map is fixed because of the frozen backbone, and is independent of a query mask. Thus we can use a gradient decomposition strategy that simplifies the gradient computation associated with the explainability method. Finally, while the resulting region-based localized embedding tokens are optimized for their specific mask, it can sometimes be desirable to also include global context. While e.g. for classification it does not matter if a bicycle is leaned to a tree or floating in the sky, such context information can be critical for referring expressions. We therefore further propose an add-on regularization loss that aligns the learned representation to the global image representation and allow to balance between global and local representations if needed.
@@
The localized embeddings produced by our method can be used in various downstream tasks as shown in Fig. 1, including region-based localized classification, region-based localized captions (as in AlphaCLIP), as well as localized image generation. In all cases, we assume a zero-shot setting and use our localized embedding tokens as a drop-in replacement, e.g. for the CLIP ViT [CLS] token. This means we compute the localized embedding token for region-based zero-shot classification and match its representation with the respective class prompts, e.g. “A photo of a dog”. We evaluate the proposed method in all those scenarios, showing improved performance compared to other methods in each domain. We summarize the contributions of our work as follows:
@@
(1) Given an image and a query mask, we learn a localized embedding at test time that captures the region characteristics within the mask in a single token. The learned can be used as a drop-in replacement for any application based on the same backbone. (2) We propose gradient decomposition to make the process computationally efficient for multiple query masks in the same image. (3) We evaluate the resulting representation on various region-based downstream tasks, showing improved results across a range of different applications.
Figure 2: MaskInversion: (Step 0): the input image is forwarded only once during the whole MaskInversion process. (Step 1): the localized embedding ($LET_{\textbf{m}}$) is initialized to the vision encoder’s [CLS] token. Then the $LET_{\textbf{m}}$ embedding is trained such that its explainability map is similar to the query mask. (Step K): after $K$ gradient descent iterations, we obtain the final localized embedding $LET_{\textbf{m}}$ that can be used for downstream task.
@@ 2 Related Work
Localized Representation Learning The task of enhancing the localized embedding of foundation models such as CLIP Radford et al. (2021) has gained increased attention recently. These models, trained on noisy image-text pairs scraped from the internet, have proven to be a rich source of supervision for learning a broad range of concepts Radford et al. (2021). However, their training methodology, which matches the global feature representation of an entire image with its corresponding caption, often falls short in the context of localized tasks. ReCLIP (Subramanian et al., 2022) uses a combination of clipping and blurring to receive a region-specific embedding and further tries to capture relations between those instances. Shtedritski et al. (Shtedritski et al., 2023) found that a red circle around an object can direct the model’s attention to that region, thus producing a ’localized’ CLS token while maintaining global information. As an extension to those works, Yang et al. Yang et al. (2024) explore different techniques for Fine-Grained Visual Prompting (FGVP), including outlining the relevant object or blurring the rest of the image (Blur Reverse Mask) and using the resulting CLIP CLS token for various downstream tasks. We find that especially the masked blurring provides a strong baseline. Another line of work, CPT (Yao et al., 2024) fine-tunes an existing language model to allow for a prompting based on different color patches. AlphaCLIP Sun et al. (2024) takes a similar approach by retraining CLIP to take an alpha mask alongside the original image as input, focusing the model’s output feature representation on the area covered by the alpha mask. However, this method requires millions of mask annotations to generalize effectively. Note that MaskInversion differs from both streams of work: from current visual prompt tuning methods, as it does not seek to change the input image directly to get a localized CLS token embedding, but instead learns a new representation for the given maks, but also from methods that rely on masked-based pertaining as MaskInversion is applied at test time and does not assume any adaptation of weights of the frozen backbone. Finally, Gal et al. (Gal et al., 2023) proposed text inversion as an idea related to capture embeddings, but for the case of learning a token that represents a certain object to be injected into a text-to-image generator. While this idea is the conceptual inspiration for this work, MaskInversion differs from this method as it captures regional properties via binary masks and respective explanation maps, while text inversion focuses on learning general object properties from multiple images.
@@
Explainability Methods The proposed MaskInversion method relies on the use of explainability methods to guide the model to focus on the desired area in the image. These methods explain model decisions by assigning a score to each image pixel representing its importance to the model’s output. Gradient-based methods, which compute explanations based on the gradient of the model’s prediction with respect to the model output, are computationally efficient and easy to understand since they are a direct function of the model’s parameters and do not rely on additional models or image modifications. They have been used successfully to identify reasoning, spurious correlation, and trustworthiness in traditional computer vision models Erhan et al. (2009); Simonyan et al. (2014); Springenberg et al. (2015); Sundararajan et al. (2017); Selvaraju et al. (2017); Smilkov et al. (2017); Kapishnikov et al. (2019). Furthermore, gradient-based methods are differentiable, making it possible to use them as an objective function. For instance, Chefer et al. (2022) uses the explainability map to supervise the model training, enforcing the model to base its classification prediction on the part of the image that contains the object, thus enhancing the model’s robustness. Similarly, Paiss et al. (2022) leverages the explainability signal to force an image generation model to utilize the entirety of the text prompt given by the user. While early explainability methods were developed for Convolutional Networks, with perhaps the most known one being GradCAM Selvaraju et al. (2017), the widespread use of ViTs has led researchers to adapt existing methods or develop new ones specifically for transformers. For instance, rollout Abnar & Zuidema (2020) combines all the attention maps via matrix multiplication to trace the flow of importance through the transformer’s layers. Chefer et al. Chefer et al. (2021) extended rollout by weighting the attention by their gradient, making the method class-specific. Recently, LeGrad Bousselham et al. (2024) proposed a gradient-based feature-attribution method specifically designed for ViT architectures. The method relies solely on the gradient of the attention maps, making it fast and easy to use. We chose LeGrad as the default explainability method used in the evaluation, but note that MaskInversion is a general method and can be used in conjunction with any differentiable explainability method.
3 Method
@@
The proposed method, coined as MaskInversion, aims to learn a localized embedding or feature vector that encapsulates an object’s characteristics within an image specified by a query mask. This embedding should not solely represent the object’s intrinsic properties but also capture the broader context of the entire image. For instance, the embedding of a mask of a cat should differ when the cat is situated in an empty field or when it is crossing a bustling road.
To achieve this, we utilize representations provided by foundation models, such as CLIP. Our approach learns a token that captures the foundation model’s feature representation on the image region specified by the mask. Hence, the foundation model remains fixed during our process.
@@
As shown in Figure 2, we start with the initialization of an embedding vector that serves as a localized embedding token of the mask. This vector is then refined through an iterative optimization process guided by an explainability map generated from the foundation model. The explainability map provides a visual indication of the areas within the image that are most influential on the initial embedding, thereby allowing for targeted refinement. The optimization process is supervised by enforcing the generated explainability map to be similar to the query mask. We can optionally use a regularization loss to ensure the mask embedding is congruent with the model’s learned manifold. Finally, we improve the computational load for this process, especially for the case of computing multiple embeddings based of different masks for the same image, via gradient decomposition.
3.1 Background/Preliminaries: Explainability Methods
@@
The proposed MaskInversion method relies on the use of explainability methods to guide the creation of the localized embedding token. Here, we give a brief introduction to explainability methods, focusing on “gradient-based” methods (e.g. GradCAMSelvaraju et al. (2017)). We let $\mathcal{F}$ denote a model that maps an input image $x\in\mathbb{R}^{3\times W\times H}$ to an output an activation $\mathcal{F}(x)=s\in\mathbb{R}$ In practice, $s$ could be derived from a classifier’s score for a particular class or the cosine similarity between image and text embeddings in a vision-language model (e.g., CLIP). For a given layer $l\in\{1,\dots,L\}$ of $\mathcal{F}$, we denote $\mathbf{A}^{l}$ the intermediate representation of the model. $\mathbf{A}^{l}$ can be intermediate features maps in the case of CNNsSelvaraju et al. (2017), intermediate tokens or attention maps in the case of Vision TransformersDosovitskiy et al. (2021). We also denote the partial derivative of the activation $s$ w.r.t $A^{l}$ as $\nabla A^{l}=\frac{\partial s}{\partial A^{l}}$.
@@
Gradient-based explainability methods can be generally formulated as combination of operations between the intermediate representation $\mathbf{A}=(A^{1},\dots,A^{L})$ and the gradients $\nabla\mathbf{A}=(\nabla A^{1},\dots,\nabla A^{L})$: and produces a 2D heatmap denoted, $E=g(\mathbf{A},\nabla\mathbf{A})\in\mathbb{R}^{W\times H}$. For instance, in GradCAM Selvaraju et al. (2017), $E$ is defined as $E(\mathbf{A},\nabla\mathbf{A})=\text{ReLU}\left(\sum_{k}\alpha_{k}\cdot\mathbf{A}^{L}_{k}\right)$, where $\alpha_{k}=\sum_{ij}\nabla A^{L}_{k,i,j}$ are the weights for the feature maps $\mathbf{A}^{L}$.
In the context of Vision Transformers (ViTs), we employ LeGrad Bousselham et al. (2024). It focuses on the attention mechanism’s role in aggregating information into the [CLS] token, which is crucial for ViTs. It considers the intermediate representations $\mathbf{A}^{l}$ to be the attention maps of the self-attention layers. For a given activation score $s$, the gradient $\nabla A^{l}$ of $s$ with respect to the attention map $A^{l}$ is computed, and a ReLU function is applied to discard negative contributions:
$\hat{E^{l}}(s)=\frac{1}{hn}\sum_{h}\sum_{i}\mathrm{ReLU}\left(\frac{\partial s}{\partial\mathbf{A}^{l}_{h,i,.}}\right).$ (1)
where $h$ is the number of heads and $n$ is the number of visual tokens. Then the explainability maps of each layers are averaged: $\bar{E}=\frac{1}{L}\sum_{l}\hat{E}^{l}(s)$. The final explainability map is then obtained by isolating the influence of the patch tokens, reshaping it into a 2D map, and applying min-max normalization to scale the scores between 0 and 1: $E=\text{norm}(\text{reshape}(\bar{E}))$. In practice, we utilize only the last attention map of the last layer to reduce computational cost.

### v2

@@ Abstract
Contrastive vision-language foundation models have achieved tremendous results in global vision-language alignment, but still show some limitations in creating representations for specific image regions. To address this problem, we propose MaskInversion, a method that leverages the feature representations of pre-trained foundation models such as CLIP to generate a context-aware embedding for a query image region specified by a mask at test time. MaskInversion starts with initializing an embedding token and compares its explainability map, derived from the pretrained model, to the query mask. The embedding token is then subsequently refined to approximate the query region by minimizing the discrepancy between its explainability map and the query mask. During this process, only the embedding vector is updated, while the underlying foundation model is kept frozen allowing to use MaskInversion with any pre-trained model. As deriving the explainability map involves computing its gradient, which can be expensive, we propose a gradient decomposition strategy that simplifies this computation. The learned region representation can be used for a broad range of tasks, including open-vocabulary class retrieval, referring expression comprehension, as well as for localized captioning and image generation. We evaluate the proposed method on all those tasks on several datasets such as PascalVOC, MSCOCO, RefCOCO, and OpenImagesV7 and show its capabilities compared to other SOTA approaches.
@@ 1 Introduction
Foundation models such as CLIP (Radford et al., 2021), pre-trained with a contrastive loss on large-scale image-text datasets, have significantly advanced vision-language understanding. However, those models focus on a global vision-language alignment in training, matching the respective text and image class ([CLS]) tokens, thus only the globally pooled information. As a result, such models often struggle with tasks requiring precise localization or the recognition of specific image regions, necessitating novel approaches to harness their full potential. In the following, we tackle the problem of generating embeddings localized to specific image regions from pretrained vision-language models. While it is possible to obtain such embeddings via naïve solutions, e.g. by processing only the cropped region, or aggregating the local token embeddings over a mask, such simple approaches often do not yield optimal results: cropping can remove important context, while token aggregation over region features might not result in a good, aligned representation as local tokens do not always correspond to the correct representation (Zhou et al., 2022).
Figure 1: MaskInversion Applications: The proposed MaskInversion method generates a localized embedding without modifying the vision encoder, thereby enabling seamless integration as a drop-in replacement for the vision encoder output across various scenarios, such as Localized Classification to classify a specific region of an image, Localized Captioning to direct the attention of an LLM to specific parts of an image, or Localized Diffusion where the embedding is used in conjunction with a diffusion model to generate variations of specific regions of images.
@@
Different approaches have been proposed to address the problem of localized vision-language tasks: ReCLIP (Subramanian et al., 2022) uses colored boxes during training to localize the alignment between vision and language. FGVP (Yang et al., 2023) employs different masking strategies to force the model to focus on the relevant object region. AlphaCLIP (Sun et al., 2024) finetunes CLIP together with an alpha channel to highlight the region of interest. Finally, RIS (Yu et al., 2023) proposes a token masking pipeline to achieve zero-shot referring image segmentation.
@@
Following this line of works, we propose MaskInversion as a method to learn a localized embedding for a query image region specified by a mask at test time, see Figure 2. MaskInversion differs from previous methods as it does not adapt the vision-language backbone, but instead leverages the explainabilty map of a frozen backbone at test-time to optimize a representation, namely a token that captures the localized embedding for a given region mask. We start by initializing this localized embedding token from the global [CLS] token produced by CLIP. This token representation is then used to compute the initial explainability map for its current representation. We then compute the difference between the explainability map and the query mask and subsequently update the token so that its representation generates an explainability map that matches the query mask. As a result, we learn a token representation specific to the image region covered by the query mask. Note that the token representation learning process is done for each mask separately. Thus, several different localized embedding tokens are created from the same image when multiple object masks are given. We can enhance the computational efficiency in this case by exploiting the fact that the derivation of the explainability map is fixed because of the frozen backbone, and is independent of a query mask. Namely, we propose a gradient decomposition strategy that simplifies the gradient computation associated with the explainability method. Finally, while the resulting localized embedding tokens are optimized for their specific mask, it can sometimes be desirable to also include global context. We therefore propose an add-on regularization loss that aligns the learned representation to the global image representation and allows to balance between global and local representations.
@@
The localized embeddings can be used in various downstream tasks, including region-based localized classification, region-based localized captions, and localized image generation (Figure 1). In all cases, we assume a zero-shot setting and use our localized embedding tokens as a drop-in replacement, e.g. for the CLIP ViT [CLS] token. This means e.g. for region-based zero-shot classification that we compute the localized embedding token and match it with the respective class prompts, e.g. “A photo of a dog”. We evaluate the proposed method in all those scenarios, showing improved performance compared to other methods in each domain.
@@
We summarize the contributions of this work as follows: (1) Given an image and a query mask, we learn a localized embedding at test time that captures the region characteristics within the mask in a single token. The learned token can be used as a drop-in replacement for any application based on the same backbone. (2) We propose gradient decomposition to make the process computationally efficient for multiple query masks in the same image. (3) We evaluate the resulting representation on various region-based downstream tasks, showing improved results across a range of different applications ranging from referring expressions to class retrieval and localized captioning.
@@ 2 Related Work
Localized Representation Learning. The task of enhancing the localized embedding of foundation models such as CLIP (Radford et al., 2021) has gained increased attention recently. Various strategies have been proposed to leverage and enhance those backbones for localized vision-language tasks. For instance, ReCLIP (Subramanian et al., 2022) uses a combination of clipping and blurring to receive a region-specific embedding and further tries to capture relations between those instances. Shtedritski et al. (Shtedritski et al., 2023) found that a red circle around an object can direct the model’s attention to that region, thus producing a ‘localized’ [CLS] token while maintaining global information. As an extension to those works, Yang et al. (Yang et al., 2023) explore different techniques for Fine-Grained Visual Prompting (FGVP), including outlining the relevant object or blurring the rest of the image (Blur Reverse Mask) and using the resulting CLIP CLS token for various downstream tasks. We find that especially the masked blurring provides a strong baseline. Another line of work, CPT (Yao et al., 2024) fine-tunes an existing language model to allow for a prompting based on different color patches. AlphaCLIP (Sun et al., 2024) takes a similar approach by retraining CLIP to take an alpha mask alongside the original image as input, focusing the model’s output feature representation on the area covered by the alpha mask. However, this method requires millions of mask annotations to generalize effectively. Note that MaskInversion differs from both streams of work. In contrast to current visual prompt tuning methods (Shtedritski et al., 2023; Yu et al., 2023; Yang et al., 2023) , it does not seek to change the input image directly to get a localized CLS token embedding, but instead learns a new representation for the given mask. In contrast to methods that rely on masked-based pretraining Sun et al. (2024); Yao et al. (2024), MaskInversion is applied at test time and does not assume any adaptation of weights of the frozen backbone. Finally, Gal et al. (Gal et al., 2023) proposed text inversion as an idea to capture embeddings in a token that represents a certain object to be injected into a text-to-image generator. While serving as inspiration for this work, MaskInversion differs from text inversion as it captures regional properties via binary masks and respective explanation maps, whereas text inversion focuses on learning general object properties from multiple images.
@@
Explainability Methods. Gradient-based methods, which compute explanations based on the gradient of the model’s prediction with respect to the model output, are computationally efficient since they are a direct function of the model’s parameters and do not rely on additional models or image modifications. They have been used successfully to identify reasoning, spurious correlation, and trustworthiness in traditional computer vision models (Erhan et al., 2009; Simonyan et al., 2014; Springenberg et al., 2015; Sundararajan et al., 2017; Selvaraju et al., 2017; Smilkov et al., 2017; Kapishnikov et al., 2019). Furthermore, gradient-based methods are differentiable, making it possible to use them as an objective function. (Chefer et al., 2022) uses the explainability map to supervise the model training, enforcing the model to base its classification prediction on the part of the image that contains the object, thus enhancing the model’s robustness. Similarly, (Paiss et al., 2022) leverages the explainability signal to force an image generation model to utilize the entirety of the text prompt given by the user. Early explainability methods were specifically developed for Convolutional Networks, e.g. GradCAM (Selvaraju et al., 2017), GradCAM++ (Chattopadhay et al., 2018) and Integrated Gradients (Sundararajan et al., 2017). However, the widespread use of Vision Transformers (ViT) has led researchers to adapt or develop methods specifically for transformers. Rollout (Abnar and Zuidema, 2020) combines all the attention maps via matrix multiplication to trace the flow of importance through the transformer’s layers. Chefer et al. (Chefer et al., 2021) extended rollout by weighting the attention by their gradient, making the method class-specific. Recently, LeGrad (Bousselham et al., 2025) proposed a gradient-based feature-attribution method specifically designed for ViT architectures relying on the gradient of the attention maps, making it fast and easy to use. We choose LeGrad as the default explainability method used in the evaluation, but note that MaskInversion is a general method and can be used with any differentiable explainability method.
Figure 2: Overview of the proposed method: Step 0: the input image is forwarded only once during the whole MaskInversion process. Step 1: the localized embedding token $LET_{\textbf{m}}$ is initialized by the vision encoder’s [CLS] token. The $LET_{\textbf{m}}$ is then trained such that its explainability map correlates to the query mask. Step K: after $K$ gradient descent steps, we obtain the final localized embedding $LET_{\textbf{m}}$ that can be used for downstream tasks.
3 Methodology
@@
The proposed method, coined as MaskInversion, aims to learn a localized embedding or feature vector that encapsulates an object’s characteristics within an image specified by a query mask.
@@
As shown in Figure 2, our method starts with the initialization of an embedding vector that serves as the localized embedding token for the mask. This vector is iteratively refined through an optimization process guided by an explainability map which highlights the image regions most influential on the embedding, enabling targeted refinement. The optimization is supervised by enforcing similarity between the generated explainability map and the query mask. Optionally, a regularization loss can be applied to enforce the mask embedding to align with the model’s learned manifold. Finally, we propose a gradient decomposition strategy to enhance computational efficiency, particularly when computing multiple embeddings for different masks on the same image.
3.1 Preliminaries: Explainability Methods
The proposed method relies on the use of explainability methods to guide the creation of the localized embedding token. Here, we give a brief introduction to explainability methods, focusing on “gradient-based” methods (e.g. GradCAM (Selvaraju et al., 2017)).
@@ 3.1 Preliminaries: Explainability Methods
Let $\mathcal{F}$ denote a model that maps an input image $x\in\mathbb{R}^{3\times W\times H}$ to an output activation $\mathcal{F}(x)=s\in\mathbb{R}$, where $s$ can be derived from a classifier’s score or the cosine similarity between image and text embeddings in a vision-language model (e.g. CLIP). We denote $\mathbf{A}^{l}$ the intermediate representation of $\mathcal{F}$ at a given layer $l\in\{1,\dots,L\}$. $\mathbf{A}^{l}$ can be intermediate features maps in the case of CNNs (Selvaraju et al., 2017), intermediate tokens or attention maps in the case of ViTs (Dosovitskiy et al., 2021). We also denote the partial derivative of the activation $s$ w.r.t $A^{l}$ as $\nabla A^{l}=\frac{\partial s}{\partial A^{l}}$.
@@
A gradient-based explainability method can be generally formulated as a combination of operations between the intermediate representation $\mathbf{A}=(A^{1},\dots,A^{L})$ and the gradients $\nabla\mathbf{A}=(\nabla A^{1},\dots,\nabla A^{L})$. It produces a 2D heatmap, denoted $E=g(\mathbf{A},\nabla\mathbf{A})\in\mathbb{R}^{W\times H}$. For instance, in GradCAM (Selvaraju et al., 2017), $E$ is defined as $E(\mathbf{A},\nabla\mathbf{A})=\text{ReLU}\left(\sum_{k}\alpha_{k}\cdot\mathbf{A}^{L}_{k}\right)$, where $\alpha_{k}=\sum_{ij}\nabla A^{L}_{k,i,j}$ are the weights for the feature maps $\mathbf{A}^{L}$. In the context of ViTs, we employ LeGrad (Bousselham et al., 2025), which considers the intermediate representations $\mathbf{A}^{l}$ to be the attention maps of the self-attention layers. For a given activation score $s$, the gradient $\nabla\mathbf{A}^{l}$ of $s$ with respect to the attention map $\mathbf{A}^{l}$ is computed, and a ReLU function is applied to discard negative contributions. While LeGrad averages the explainability maps of several layers, we here utilize only the attention map of the last layer to reduce computational cost.